OpenStudy (anonymous):

Factor completely 125x^3-64

OpenStudy (anonymous):

Difference of cubes. a3 + b3 = (a + b)(a2 – ab + b2) a3 – b3 = (a – b)(a2 + ab + b2) (5x-4)(25x^2+20x+16)